OpenStudy (anonymous):

Solve by using the perfect squares method. x2 – 12x + 36 = 0 A. 18 B. –18 C. 6 D. –6

OpenStudy (anonymous):

c x2-6x-6x+36=0 x(x-6)-6(x-6)=0 x-6=0 x=6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome :)